OpenStudy (animalover02):

I WILL MEDAL & FAN!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Elana bought 320 tropical fish for a museum display . she bought 7 times a s many triggerfish as parrotfish. How many of each type of fish did she buy? A. 40 parrotfish , 280 triggerfish B. 280 parrotfish , 40 triggerfish C. 20 parrotfish , 140 triggerfish D. 140 parrotfish , 200 triggerfish

OpenStudy (animalover02):

@M.D.Saxon

OpenStudy (m.d.saxon):

its A

OpenStudy (animalover02):

thanks!

OpenStudy (animalover02):

@shaik0124 do you also think its A?

OpenStudy (shaik0124):

it's like assume the number of parrotfish as x and in the question it is given as rigger fish as 7 times parrotfish i.e,7x now 7x+x=320 8x=320 x=40(parrotfish) 7x=7*40=280(triggerfish)